Nitrogen gas reacts with hydrogen gas to form ammonia gas. Given that the standard entropies of nitrogen, hydrogen and ammonia gases are 192, 131 and 193 JK-¹mol-¹ respectively, calculate the standard entropy change for the reaction


1
Expert's answer
2021-03-25T04:10:32-0400

1/2 N2 + 3/2 H2 > NH3

∆ S ° = ∆S° (product) - ∆S° (Reactant)

∆S° = 192- (1/2 ×191.5 + 3/2 × 130.5)

∆S° = 192.6 - (95.75 + 195.75)

∆S° = 192.6 - 291.5 = -98.9 J/mol -k

Therefore ∆S° of the formation of Ammonia is -98.9 J/mol -k